Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b.
(7,9); x+7y=2
The equation of the line is?
(Show the equation. Show answer in slope-intercept form. Use integers or fractions for any numbers in the equation. Simplify the answer.)

Answer by stanbon(75887)

Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b.
(7,9); x+7y=2
----
Given-line DATA::
Form y = -x/7 + 2/7
Slope = -1/7
----------------------
Desired-line::
Form: y = mx + b
Solve for "m" and "b"
m = -1/7
----
Solve for "b":
9 = (-1/7)7 + b
b = 10
--------
Equation:
y = (-1/7)x + 10
--------------------------
Cheers,
Stan H.
----------------------